# A new rotating black hole from the Newman-Janis algorithm

### v1

@@ Abstract
We present a new rotating black hole solution to the Einstein equations as an extension of the Kerr spacetime. To derive this solution, we use the Newman-Janis algorithm as a mathematical tool that reduces a general rotating metric into a tractable form by applying simple physical requirements. Interestingly, the solution we found may not be uniquely characterized by asymptotic parameters such as mass, angular momentum, and charge, thereby challenging the no-hair theorem. We also analyze in detail how this additional characteristics (“hair”) affects the thermodynamic properties of the black hole.
@@ 1 Introduction
This naturally raises the question: How can we derive solutions for rotating black holes that coexist with surrounding matter fields? The first major extension in this direction was the Kerr-Newman solution [9], which describes a rotating, charged black hole in Einstein-Maxwell theory . This solution was derived using the Newman-Janis (NJ) algorithm [10], a method that has been widely studied and debated [11, 12, 13]. While the underlying reason how the algorithm works remains unclear, its empirical success makes it a valuable tool for deriving rotating black hole solutions. Recent research has focused on extending this algorithm to incorporate additional matter fields [14, 15, 16, 17]. There is also considerable interest in which types of fields can coexist with rotating black holes while maintaining physical consistency [18, 19, 17]. In various theories of gravity, efforts have been made new rotating black hole solutions in different spacetime dimensions [20, 21, 22, 23, 24, 25, 26, 27].
@@
To study the matter fields coexisting with a rotating black hole, it is crucial to analyze the components of the stress-energy tensor and understand the equation of state, which determines the nature of the matter [28]. Such an analysis provides valuable insights into the behavior of the matter fields surrounding the rotating black holes. For the electrically charged Kerr-Newman black hole, the stress-energy tensor reveals that the energy density is related to the radial pressure by the equation of state $\varepsilon=-p_{r}$. Interestingly, Maxwell’s field having this property does not behave as a perfect fluid. When considering more general matter fields, two main approaches emerge: First is to find a rotating black hole solution that coexists with matter of the perfect fluid type [29, 30]. The second involves extending the Kerr-(Newman) black hole to the rotating black hole with matter fields having the property $\varepsilon=-p_{r}$, such as electric fields [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41].
@@
The no-hair theorem is well-established for the electrovacuum solution of Einstein-Maxwell theory [64], which states that black holes are characterized solely by their mass, charge, and angular momentum as primary hairs. However, in Einstein-Maxwell theory with additional matter fields beyond ordinary matter or in extensions of general relativity, the applicability of the no-hair theorem is not necessarily evident. When fields other than the electromagnetic field are present, additional hairs may emerge. In the Einstein-scalar theory with minimal coupling, Bekenstein’s no-hair theorem holds under the assumption that the energy density of the scalar field outside the black hole horizon is non-negative [42, 43]. Thus, the theorem can be circumvented if this assumption is violated [44, 45, 46, 47, 48, 49, 50, 51]. Further investigation is required to determine whether such hair should be classified as primary or secondary [52, 53, 54].
@@
In this work, we explore the Newman-Janis algorithm in a general setting and analyze its role in constructing rotating black hole geometries that coexist with matter fields. Our approach extends beyond the standard Kerr-Newman solution, allowing for more general modifications allowing the new kind of hairs while maintaining physical consistency. We investigate the structure of these solutions and discuss their implications for black hole physics and astrophysical observations.
The paper is organized as follows: In Secs. 2 and 3, we investigate the mathematical structure of the NJ algorithm more closely and derive a new solution, which is a generalization of the Kerr black hole coexisting with a matter field. Sec. 4 is dedicated to analyzing the properties of this new solution in depth. In Sec. 5, we explore the implications of the new solution for black hole thermodynamics, with a focus on how the presence of the new matter field affects thermodynamic properties. Finally, we summarize our findings and discuss their broader implications in the concluding section.
2 Mathematical structure of the Newman-Janis algorithm
In this section, we revisit and analyze the mathematical structure of the NJ algorithm in a general setting. The goal is to provide a detailed mathematical analysis of the original NJ algorithm to derive a new solution corresponding to a rotating black hole geometry. To ensure that the resulting metric describes physical spacetime, it is necessary to verify that the metric satisfies the Einstein equation and derive the corresponding energy-momentum tensor, which will be done in the next section. Throughout this analysis, we assume that the spacetime geometry is asymptotically flat.
As the base geometry, we consider a static, spherically symmetric metric of the form,
@@ 2 Mathematical structure of the Newman-Janis algorithm
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta,d\phi^{2}$ represents the angular part of the metric. When $\sigma(r)=r^{2}$, the radial coordinate $r$ corresponds to the areal radius. However, If $\sigma(r)\neq r^{2}$, the areal radius becomes $\sqrt{\sigma(r)}$. At this stage, we do not presume any specific forms for the functions $f(r)$, $g(r)$, and $\sigma(r)$ aside from the asymptotic flatness condition, which requires $f(r)\,,g(r)\to 1$ and $\sigma(r)\to r^{2}$ as $r\to\infty$.
Let us apply the Newman-Janis algorithm step by step.
  1. 1.
    The first step involves expressing the metric in null coordinates and constructing the corresponding tetrad. We introduce the null coordinate $u$ via the transformation, $du=dt-\frac{dr}{\sqrt{f(r)g(r)}}$, which allows us to rewrite the metric (1) in the Eddington-Finkelstein form:
    $ds^{2}=-f(r)du^{2}-2\sqrt{\frac{f(r)}{g(r)}}dudr+\sigma(r)d\Omega^{2}\,.$ (2)
    Next, we construct a set of tetrads $\{l_{\mu},n_{\mu},m_{\mu},\bar{m}_{\mu}\}$ satisfying the pseudo-orthogonality relations: $l^{\mu}n_{\mu}=-1$ and $m^{\mu}\bar{m}_{\mu}=1$, with the metric expressed as:
    $g_{\mu\nu}=-l_{\mu}n_{\nu}-n_{\mu}l_{\nu}+m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu}\,.$ (3)
    The components of the tetrad are then given by
    $l_{\mu}=\delta_{\mu}^{0}\,,\quad n_{\mu}=\frac{f(r)}{2}\delta_{\mu}^{0}+\sqrt{\frac{f(r)}{g(r)}}\delta_{\mu}^{1}\,,$
    $m_{\mu}=\frac{\sqrt{\sigma(r)}}{\sqrt{2}}(\delta_{\mu}^{2}+i\sin\theta\,\delta_{\mu}^{3})\,,\quad\bar{m}_{\mu}=\frac{\sqrt{\sigma(r)}}{\sqrt{2}}(\delta_{\mu}^{2}-i\sin\theta\,\delta_{\mu}^{3})\,.$ (4)
    It is straightforward to verify that this tetrad satisfies the required conditions. The tetrad components with upper indices can be computed as:
    $l^{\mu}=-\sqrt{\frac{g(r)}{f(r)}}\delta^{\mu}_{1}\,,\quad n^{\mu}=\left(-\delta^{\mu}_{0}+\frac{\sqrt{f(r)g(r)}}{2}\delta^{\mu}_{1}\right)\,,\quad m^{\mu}=\frac{1}{\sqrt{2\sigma(r)}}\left(\delta^{\mu}_{2}+\frac{i}{\sin\theta}\delta^{\mu}_{3}\right)\,.$ (5)
    This tetrad construction forms the foundation for the application of the NJ algorithm to derive the rotating solution.
  2. 2.
    The second step involves performing a complex coordinate transform from $(u,r)\to(u^{\prime},r^{\prime})$  [^3] In the original form, $p(x)=q(x)=x=\cos\theta$. More general transform could be assumed as in Ref. [11]. as follows:
    $u^{\prime}=u-iap(x),\qquad r^{\prime}=r+iaq(x),\qquad x=\cos\theta\,,$ (6)
    where $a$ denotes the rotation parameter. The angle dependent functions $p(x)$ and $q(x)$ denote distinct coordinates transforms for $u$ and $r$. Notably, we do not assume $p(x)=q(x)$ in general. Later in this work, we find $p(x)=q(x)$ for the Schwarzschild case, dictated by physical requirements rather than purely mathematical reasons.
    In this new coordinates, $x^{\mu}=(u^{\prime},r^{\prime},\theta,\phi)$, the tetrad vectors are expressed as:
    $l^{\mu}$ $=$ $-\sqrt{\frac{G(r,\theta)}{F(r,\theta)}}\delta^{\mu}_{1},$
    $n^{\mu}$ $=$ $-\delta^{\mu}_{0}+\frac{\sqrt{F(r,\theta)G(r,\theta)}}{2}\delta^{\mu}_{1}\,,$
    $m^{\mu}$ $=$ $\frac{1}{\sqrt{2\Sigma(r,\theta)}}\left[ia\sin\theta({\dot{p}}(x)\,\delta^{\mu}_{0}-{\dot{q}}(x)\delta^{\mu}_{1})+\delta^{\mu}_{2}+\frac{i}{\sin\theta}\,\delta^{\mu}_{3}\right]\,,$ (7)
    where ${\dot{p}}(x)\equiv(dp(x)/dx)$ and ${\dot{q}}\equiv(dq(x)/dx)$. Here, $F(r,\theta)$, $G(r,\theta)$, and $\Sigma(r,\theta)$ represent generalizations of the functions $f(r)$, $g(r)$, and $\sigma(r)$ in the presence of the rotation parameter $a$, respectively. In performing this generalization, one usually adopt the prescriptions
    $r^{2}\to(r+iaq)(r-iaq),\qquad\frac{2}{r}\to\frac{1}{r+iaq}+\frac{1}{r-iaq}\,.$ (8)
    However, as noted in Ref. [11], ambiguities arise in applying these prescriptions, particularly in the second case. These ambiguities cannot be fully resolved by mathematical reasoning along and are addressed using physical arguments in the next section.
    Using the tetrad, we reconstruct the inverse metric from the relation $-l^{\mu}n^{\nu}-n^{\mu}l^{\nu}+m^{\mu}\bar{m}^{\nu}+\bar{m}^{\mu}m^{\nu}=g^{\mu\nu}$. Inverting the inverse metric yields the line elements in the new coordinates $(u^{\prime},r^{\prime},\theta,\phi)$:
    $ds^{2}$ $=$ $-F(r,\theta)d{u}^{2}-2\sqrt{\frac{F(r,\theta)}{G(r,\theta)}}dudr-2a\left({\dot{q}}(x)\sqrt{\frac{F(r,\theta)}{G(r,\theta)}}-{\dot{p}}(x)F(r,\theta)\right)\sin^{2}\theta dud\phi$ (9)
    $+2a\sqrt{\frac{F(r,\theta)}{G(r,\theta)}}{\dot{p}}(x)\sin^{2}\theta\,drd\phi+\Sigma(r,\theta)d\theta^{2}$
    $+\left(\Sigma(r,\theta)+{\dot{p}}(x)\left(2{\dot{q}}(x)\sqrt{\frac{F(r,\theta)}{G(r,\theta)}}-{\dot{p}}(x)F(r,\theta)\right)a^{2}\sin^{2}\theta\right)\sin^{2}\theta\,d\phi^{2}\,.$
    For simplicity, we omit the primes for the coordinates.
  3. 3.
    The next step in the NJ algorithm is to transform the metric into a rotating form by removing the off-diagonal components $dtdr$ and $drd\phi$. To achieve this, we introduce the following coordinate transformations:
    $du=dt-\frac{\Gamma(r)dr}{\Delta(r)},\qquad d\phi=d\psi-\frac{adr}{\Delta(r)}\,.$ (10)
    Since $u$ and $\phi$ are coordinates bases, the transformations in equation (10) must satisfy the condition of integrability. This implies that the functions $\Gamma(r)$ and $\Delta(r)$ depend only on $r$ and are independent of $\theta$.
    This dependency is significant as it greatly simplifies the process of solving the Einstein equation in a rotating system. The integrability condition also ensures that the transformed metric components $g_{tr}$ and $g_{r\psi}$ vanish, as required to describe a stationary rotating solution with a symmetry axis aligned along the $z$-direction.
    To enforce these constraints, the form of $F(r,\theta)$ and $G(r,\theta)$ are determined in terms of $\Delta(r)$ and $\Gamma(r)$ as follows:
    $F(r,\theta)=\frac{\Sigma(r,\theta)\big{(}\Delta(r)-a^{2}{\dot{q}}(x)\sin^{2}\theta\big{)}}{{\dot{q}}(x)\left(\Gamma(r)-a^{2}{\dot{p}}(x)\sin^{2}\theta\right)^{2}}\,,~{}~{}G(r,\theta)=\frac{{\dot{q}}(r)}{\Sigma(r,\theta)}\left[\Delta(r)-{\dot{q}}(x)a^{2}\sin^{2}\theta\right]\,.$ (11)
    With these forms, the metric is transformed into a Boyer-Lindquist form [56]
    $ds^{2}$ $=$ $-F(r,\theta)dt^{2}-2a\left({\dot{q}}(x)\sqrt{\frac{F(r,\theta)}{G(r,\theta)}}-{\dot{p}}(x)F(r,\theta)\right)\sin^{2}\theta dtd\psi+\Sigma(r,\theta)d\theta^{2}$ (12)
    $+\frac{1}{{\dot{q}}(x)}\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}$
    $+\left[\Sigma(r,\theta)+a^{2}\sin^{2}\theta\,{\dot{p}}(x)\left(2{\dot{q}}(x)\sqrt{\frac{F(r,\theta)}{G(r,\theta)}}-{\dot{p}}(x)F(r,\theta)\right)\right]\sin^{2}\theta\,d\psi^{2}\,.$
    At this stage, the NJ algorithm formally concludes, providing the general rotating metric. Further analysis, such as checking the Einstein equations and determining the physical implications, is required to finalize the solution.
As demonstrated in the next section, the Kerr and the Kerr-Newman black holes can be derived using the three steps by choosing $p(x)=q(x)=x$ and applying the prescription (8). With the specific choice, the Einstein equation is automatically satisfied. However, in general cases, the metric (12) must independently satisfy the Einstein equations without relying solely on the prescription (8). This opens the possibility of exploring different prescriptions or alternative functional forms for $p(x)$ and $q(x)$, provided the resulting metric (12) remains consistent with the Einstein equation. In the next section, we investigate whether such alternative choices are feasible and physically meaningful.
The determinant of the metric (12) is given by
$\det g_{\mu\nu}=-\frac{F(r,\theta)\Sigma^{2}(r,\theta)}{G(r,\theta)}\sin^{2}\theta=-\left(\frac{\Sigma^{2}(r,\theta)\sin\theta}{{\dot{q}}(x)(\Gamma(r)-a^{2}{\dot{p}}(x)\sin^{2}\theta)}\right)^{2}\,.$
Note that the coefficient of $dr^{2}$ becomes $g_{rr}(r,\theta)=\frac{1}{{\dot{q}}(x)}\frac{\Sigma(r,\theta)}{\Delta(r)}$ Here, $\Sigma(r,\theta)$ and $\Delta(r)$ are positive-definite asymptotically. To avoid negative values for $g_{rr}$ in the asymptotic region and to prevent an unwanted singularity in $\det g_{\mu\nu}$, we require
${\dot{q}}(x)>0,\quad\Gamma(r)>|a^{2}{\dot{p}}(x)|\quad\mbox{ for all }|x|\leq 1\,.$ (13)
A typical choice is ${\dot{q}}(x)=1$ as the $\theta$-dependence in the $g_{rr}$ term, except for $\Sigma(r,\theta)$, could disrupt the axial symmetry of the system.
We now rewrite this metric to the form of an orthonormal, non-coordinate basis frame. Additionally, we replace the angle $\psi\to\phi$, yielding
$ds^{2}$ $=$ $-\frac{\Sigma(r,\theta)\Delta(r)}{{\dot{q}}(x)\left(\Gamma(r)-{\dot{p}}(x)a^{2}\sin^{2}\theta\right)^{2}}\left(dt-{\dot{p}}(x)a\sin^{2}\theta d\phi\right)^{2}$ (14)
$+\Sigma(r,\theta)d\theta^{2}+\frac{1}{{\dot{q}}(x)}\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}$
$+\frac{\Sigma(r,\theta)\sin^{2}\theta}{\left(\Gamma(r)-{\dot{p}}(x)a^{2}\sin^{2}\theta\right)^{2}}(\Gamma(r)d\phi-adt)^{2}\,.$
@@
The metric (14) is a mathematical generalization of the general spherically symmetric metric (1) by the Newman-Janis algorithm. In this metric, there are five unknown functions to be determined. Of these, only $\Sigma(r,\theta)$, which generalizes $\sigma(r)$, depends on both $r$ and $\theta$. The coordinate transformation functions $p(x)$ and $q(x)$ are functions of $\theta$ only, while $\Gamma(r)$ and $\Delta(r)$ depend only on $r$. The physical context should specify the appropriate choices for these functions. Rewriting the metric in this form is essential, as the Einstein tensors in this basis, for geometries like the Kerr, Kerr-Newman black holes, take diagonal forms.
Note that the functions $F(r,\theta)$ and $G(r,\theta)$ are the generalization of $f(r)$ and $g(r)$. Therefore, in the non-rotating limit, we have
$g(r)\equiv\lim_{a\to 0}G(r,\theta)=\lim_{a\to 0}\frac{{\dot{q}}(x)\Delta(r)}{\Sigma(r,\theta)},\qquad f(r)\equiv\lim_{a\to 0}F(r,\theta)=\lim_{a\to 0}\frac{\Delta(r)\Sigma(r,\theta)}{{\dot{q}}(x)\Gamma^{2}}\,.$ (15)
@@
Noting $\Sigma(r,\theta)\to\sigma(r)$, ${\dot{q}}(x)\to 1$ in the $a\to 0$ limit, this relation constrains the functional forms for $\Delta(r)$ and $\Gamma(r)$. Since $\sigma(r)$ is independent of $\theta$, we can conclude that ${\dot{q}}(x)$ must be independent of $x$ in the non-rotating limit.
@@ 3 Rotating geometries from the static spherically symmetric black hole
In this section, we demonstrate how to generate rotating geometries starting from a static one by applying the Newman-Janis (NJ) algorithm, as outlined in the previous section. We begin with the Schwarzschild metric and use the method to construct not only the Kerr geometry but also several new rotating geometries.
The Schwarzschild metric, which describes a static, spherically symmetric black hole, is given by
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{dr^{2}}{1-2M/r}+r^{2}d\Omega^{2},\qquad r\geq 0\,.$ (16)
For this metric, the associated functions are:
$f(r)=g(r)=1-\frac{2M}{r},\qquad\sigma(r)=r^{2}\,.$
Since the Schwarzschild black hole is a vacuum solution, the Einstein tensor that corresponds to this geometry vanishes, satisfying the Einstein field equations in the absence of any matter. Now, we proceed to develop a rotating geometry step by step using the NJ algorithm.
3.1 The Kerr geometry
To illuminate the procedure, we step through the derivation of the Kerr geometry using the Newman-Janis (NJ) algorithm, starting from the Schwarzschild metric. Here’s the process broken down:
  1. 1.
    We begin by choosing the coordinate transformation functions. For simplicity, we set $q(x)=x$, which satisfies the constraint in Eq. (13).
  2. 2.
    The next step is to define the function $\Sigma(r,\theta)$ starting from $\sigma(r)=r^{2}$. Following the prescription (8), the generalization to $\Sigma(r,\theta)$ appears unambiguous,
    $\Sigma(r,\theta)=(r+iaq)(r-iaq)=r^{2}+a^{2}q^{2}=r^{2}+a^{2}\cos^{2}\theta\,.$ (17)
  3. 3.
    On the other hand, the generalization from $f(r)$ and $g(r)$ to $F(r,\theta)$ and $G(r,\theta)$ requires caution. The reason is that the functions $F(r,\theta)$ and $G(r,\theta)$ should be defined through the functions $\Delta(r)$ and $\Gamma(r)$ indirectly from Eq. (11). From the non-rotating limit in Eq. (15), we know
    $\lim_{a\to 0}\Delta(r)=r^{2}g(r)=r^{2}-2Mr\,,\qquad\lim_{a\to 0}\Gamma(r)=r^{2}\,.$ (18)
    In fact, any functions $F(r,\theta)$ and $G(r,\theta)$ satisfying the constraints are acceptable if the result is physically sound. For the case of the Kerr geometry, we follow the prescription (8) for $G$ and using Eq. (11), we have
    $g(r)$ $\to G(r,\theta)=\frac{\Sigma(r,\theta)-2Mr}{\Sigma(r,\theta)}=\frac{\Delta(r)-a^{2}\sin^{2}\theta}{\Sigma(r,\theta)}.$ (19)
    This leads to the following expression for $\Delta(r)$:
    $\Delta(r)$ $=$ $\Sigma(r,\theta)+a^{2}\sin^{2}\theta-2Mr=r^{2}+a^{2}-2Mr\,.$ (20)
    Note that this expression for $\Delta(r)$ is consistent with the non-rotating limit (18). We also follow the prescription (8) for $F(r,\theta)$ and get
    $f(r)\to F(r,\theta)=1-\frac{2Mr}{\Sigma(r,\theta)}\,.$ (21)
    Equating this form with that in Eq. (11), we get
    $\Gamma(r)-a^{2}{\dot{p}}(x)\sin^{2}\theta=\Sigma(r,\theta)=r^{2}+a^{2}\cos^{2}\theta\quad\rightarrow\quad\Gamma(r)=r^{2}+a^{2},\qquad{\dot{p}}(x)=1.$ (22)
    Here, we have used the fact that $\Gamma(r)$ is a function of $r$ only. Note that the function $p(x)=x$ is fixed at this step because of the prescription (8).
    Substituting the functions $F(r,\theta)$ and $G(r,\theta)$ into the general form of the metric, we obtain the Kerr geometry in the Boyer-Lindquist coordinates,
    $ds^{2}$ $=$ $-\frac{\Delta(r)-a^{2}\sin^{2}\theta}{\Sigma(r,\theta)}dt^{2}-\frac{2(r^{2}+a^{2}-\Delta(r))}{\Sigma(r,\theta)}a\sin^{2}\theta dtd\phi+\Sigma(r,\theta)d\theta^{2}$ (23)
    $+\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}+\left[-\frac{\Delta(r)}{\Sigma(r,\theta)}a^{2}\sin^{2}\theta+\frac{(r^{2}+a^{2})^{2}}{\Sigma(r,\theta)}\right]\sin^{2}\theta\,d\phi^{2}$
    $=$ $-\frac{\Delta(r)}{\Sigma(r,\theta)}\left(dt-a\sin^{2}\theta d\phi\right)^{2}+\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}$
    $+\Sigma(r,\theta)d\theta^{2}+\frac{\sin^{2}\theta}{\Sigma(r,\theta)}\left((r^{2}+a^{2})d\phi-adt\right)^{2}\,.$
  4. 4.
    Finally, the Einstein tensor for this geometry is calculated and found to vanish, as expected for a vacuum solution. However, it is important to be cautious when considering the physical implications of this result. The final step in applying the NJ algorithm often involves ensuring that the geometry satisfies physical conditions, such as regularity at the event horizon. We will present this aspect in the next subsection.
The first three steps are merely a mathematical extension of the Schwarzschild geometry using a complex coordinate transformation. The Kerr geometry emerges solely from this mathematical extension. However, the final step, which involves checking its physical implications, is generally necessary.
3.2 Generalization of the Kerr geometry
In this subsection, we develop new rotating solutions, which are a generalization of the Kerr geometry. We aim to find new solutions by modifying the Kerr geometry minimally.
  1. 1.
    To make a minimal change from the Kerr geometry, we choose the angle-dependent coordinates transformation functions, $q(x)=x=p(x)$. For $\Sigma(r,\theta)$, we follow the same form as the Kerr’s in Eq. (17).
  2. 2.
    In determining $F(r,\theta)$ and $G(r,\theta)$, we do not follow the standard prescription. The only physical requirement which should be maintained is the limiting condition in Eq. (18). Because the functions still be determined from $\Delta(r)$ and $\Gamma(r)$ as given in Eq. (11). Now, we choose the following general form:
@@ 3.2 Generalization of the Kerr geometry
    where $v(r)$ and $\tilde{A}(r)$ denote the deviations from the Kerr geometry. Now, the metric (14) becomes
@@
    $+\frac{\Sigma(r,\theta)\sin^{2}\theta}{\left(\Sigma(r,\theta)+\tilde{A}(r)\right)^{2}}\left(\Gamma(r)d\phi-adt\right)^{2}\,.$
  3. 3.
    In this step, we require the condition that the stress tensor does not have the off-diagonal [^4] Note however that this non-vanishing value does not imply the existence of shear viscosity between the two directions. One can find an orthonormal basis which makes this value vanish by rotating $r$ and $\theta$ directions. $r\theta$ component, $T_{\hat{r}\hat{\theta}}=0$. This is required for fluids undergoing only a rotational motion about a fixed axis (the $z$-axis here). This leads to $R_{\hat{r}\hat{\theta}}=0$ [57]. To calculate the Einstein tensor, we take the orthonormal non-coordinate basis from the metric (25):
@@
    $\omega^{\hat{\theta}}=\sqrt{\Sigma(r,\theta)}d\theta,\quad\omega^{\hat{\phi}}=\frac{\sqrt{\Sigma(r,\theta)}\sin\theta}{\Sigma(r,\theta)+\tilde{A}(r)}\left[\Gamma(r)d\phi-adt\right]\,.$ (26)
    Starting from this basis, we construct the connection 1-form through $d\omega^{\hat{a}}=-\omega^{\hat{a}}_{~{}\hat{b}}\wedge\omega^{\hat{b}}$ and find the curvature 2-form, $\mathfrak{R}^{\hat{a}}_{~{}\hat{b}}=d\omega^{\hat{a}}_{~{}\hat{b}}+\omega^{\hat{a}}_{~{}\hat{c}}\wedge\omega^{\hat{c}}_{~{}\hat{b}}.$ From the coefficients of this curvature 2-forms, we read off the curvature components. After a lengthy calculation, one may find that the $r\theta$ component of the Ricci tensor is given by
@@
    Notice that if this tensor does not vanish, it presents the non-vanishing off-diagonal stress tensor $T_{\hat{r}\hat{\theta}}$. Therefore, we choose $R_{\hat{r}\hat{\theta}}=0$, which leads

### v2

# Dressing rotating black holes with anisotropic matter

@@ Abstract
We present a new rotating black hole solution to the Einstein equations as an extension of the Kerr spacetime. To derive this solution, we use the Newman-Janis algorithm and results from the non-complexification procedure. Interestingly, the solution we find may not be uniquely characterized by asymptotic parameters such as mass, angular momentum, and charge, thereby challenging the no-hair theorem. We also analyze in detail how this additional characteristics (“hair”) affects the thermodynamic properties of the black hole.
@@ 1 Introduction
This naturally raises the question: How can we derive solutions for rotating black holes that coexist with surrounding matter fields? The first major extension in this direction was the Kerr-Newman solution [9], which describes a rotating, charged black hole in Einstein-Maxwell theory . This solution was derived using the Newman-Janis (NJ) algorithm [10], a method that has been widely studied and debated [11, 12, 13, 14, 15, 16, 17]. While the underlying reason how the algorithm works remains unclear, its empirical success makes it a valuable tool for deriving rotating black hole solutions. Recent research has focused on extending this algorithm to incorporate additional matter fields [18, 19, 20, 21]. There is also considerable interest in which types of fields can coexist with rotating black holes while maintaining physical consistency [22, 23, 21]. In various theories of gravity, efforts have been made new rotating black hole solutions in different spacetime dimensions [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
To study the matter fields coexisting with a rotating black hole, it is crucial to analyze the components of the stress-energy tensor and understand the equation of state, which determines the nature of the matter [39]. Such an analysis provides valuable insights into the behavior of the matter fields surrounding the rotating black holes. For the electrically charged Kerr-Newman black hole, the stress-energy tensor reveals that the energy density is related to the radial pressure by the equation of state $\varepsilon=-p_{r}$. Interestingly, Maxwell’s field having this property does not behave as a perfect fluid. When considering more general matter fields, two main approaches emerge: First is to find a rotating black hole solution that coexists with matter of the perfect fluid type [40, 41]. The second involves extending the Kerr-(Newman) black hole to the rotating black hole with matter fields having the property $\varepsilon=-p_{r}$, such as electric fields [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 18, 19, 20, 21].
@@
The no-hair theorem is well-established for the electrovacuum solution of Einstein-Maxwell theory [82], which states that black holes are characterized solely by their mass, charge, and angular momentum as primary hairs. However, in Einstein-Maxwell theory with additional matter fields beyond ordinary matter or in extensions of general relativity, the applicability of the no-hair theorem is not necessarily evident. When fields other than the electromagnetic field are present, additional hairs may emerge. In the Einstein-scalar theory with minimal coupling, Bekenstein’s no-hair theorem holds under the assumption that the energy density of the scalar field outside the black hole horizon is non-negative [54, 55]. Thus, the theorem can be circumvented if this assumption is violated [56, 57, 58, 59, 60, 62, 63, 64, 65, 66, 67]. Further investigation is required to determine whether such hair should be classified as primary or secondary [68, 69, 70].
@@
In this work, we use results of the “non-complexification procedure” [13], a developments of the Newman-Janis algorithm, as a method for separation of variables for the Einstein equation to explore the structure of axially symmetric, rotating solutions. Let us briefly summarize the strategy to get the rotating solution in this schime. As the base geometry, we consider a static, spherically symmetric metric of the form,
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ represents the angular part of the metric. A black hole solutions in mind, we set $\sigma(r)=r^{2}$. At this stage, we do not presume any specific forms for the functions $f(r)$ and $g(r)$ aside from the asymptotic flatness condition, which requires $f(r)\,,g(r)\to 1$ as $r\to\infty$.
The resulting rotating geometries can be written as the form
$ds^{2}$ $=$ $-\frac{\Sigma\Delta}{\left(\Gamma-a^{2}\sin^{2}\theta\right)^{2}}\left(dt-a\sin^{2}\theta d\phi\right)^{2}+\Sigma d\theta^{2}+\frac{\Sigma}{\Delta}dr^{2}$ (2)
$+\frac{\Sigma\sin^{2}\theta}{\left(\Gamma-a^{2}\sin^{2}\theta\right)^{2}}(adt-\Gamma d\phi)^{2}\,,$
where $\Sigma(r,\theta)$, $\Delta(r)$, and $\Gamma(r)$ will be shown in Sec. 2. By writing the metric function in this way, the orthonormal basis is easy to read. The components of the energy-momentum tensor are diagonalized on an orthonormal frame [72, 73].
@@
In this metric, there are three unknown functions to be determined. Of these, only $\Sigma(r,\theta)$, which generalizes $\sigma(r)$, depends both on $r$ and $\theta$. The coordinate transformation functions $\Gamma(r)$ and $\Delta(r)$ depend only on $r$. The physical context should specify the appropriate choices for these functions. The functions are determined through physical requirements and the Einstein equations with the constraints in the non-rotating limit,
$g(r)\equiv\lim_{a\to 0}\frac{\Delta(r)}{\Sigma(r,\theta)},\qquad f(r)\equiv\lim_{a\to 0}\frac{\Delta(r)\Sigma(r,\theta)}{\Gamma^{2}}\,.$ (3)
@@
Noting $\Sigma(r,\theta)\to\sigma(r)$ in the $a\to 0$ limit, this relation constrains the functional forms for $\Delta(r)$ and $\Gamma(r)$.
In this work, we regard the metric (2) as the starting one to analyze the Einstein equations and get the orthonormal frame. Therefore, we does not assume the predefined static limit (3) but try to get the solution from the general physical requirements. We investigate the structure of these solutions and discuss their implications for black hole physics.
The paper is organized as follows: In Sec. 2, we derive a new rotating solution of the Einstein equation, which is a generalization of the Kerr black hole coexisting with a matter field. Sec. 3 is dedicated to analyzing the properties of this new solution in depth. In Sec. 4, we explore the implications of the new solution for black hole thermodynamics, with a focus on how the presence of the new matter field affects thermodynamic properties. Finally, we summarize our findings and discuss their broader implications in the concluding section.
@@ 2 Rotating geometries from the static spherically symmetric black hole
We develop new rotating solutions, which are a generalization of the Kerr geometry. We aim to find new solutions by modifying the Kerr geometry minimally. Thus, we follow the same form as the Kerr’s for $\Sigma(r,\theta)$:
$\Sigma(r,\theta)=r^{2}+a^{2}\cos^{2}\theta.$ (4)
The metric functions are determined from $\Delta(r)$ and $\Gamma(r)$ as given in Eq. (2), which are prescribed through Einstein equation. Therefore, we choose the following general form:
@@
where $v(r)$ and $\tilde{A}(r)$ denote the deviations from the Kerr geometry. The metric (2) becomes
@@
$+\frac{\Sigma(r,\theta)\sin^{2}\theta}{\left(\Sigma(r,\theta)+\tilde{A}(r)\right)^{2}}\left(adt-\Gamma(r)d\phi\right)^{2}\,.$
Then, we require the condition that the stress tensor does not have the off-diagonal [^3] Note however that this non-vanishing value does not imply the existence of shear viscosity between the two directions. One could find the orthonormal basis which makes this value vanish by rotating $r$ and $\theta$ directions. $r\theta$ component, $T_{\hat{r}\hat{\theta}}=0$. This is required for fluids undergoing only a rotational motion about a fixed axis (the $z$-axis here), which leads to $R_{\hat{r}\hat{\theta}}=0$ [73]. To calculate the Einstein tensor, we take the orthonormal non-coordinate basis from the metric (6):
@@
$\omega^{\hat{\theta}}=\sqrt{\Sigma(r,\theta)}d\theta,\quad\omega^{\hat{\phi}}=\frac{\sqrt{\Sigma(r,\theta)}\sin\theta}{\Sigma(r,\theta)+\tilde{A}(r)}\left[adt-\Gamma(r)d\phi\right]\,.$ (7)
Starting from this basis, we construct the connection 1-form through $d\omega^{\hat{a}}=-\omega^{\hat{a}}_{~{}\hat{b}}\wedge\omega^{\hat{b}}$ and find the curvature 2-form, $\mathfrak{R}^{\hat{a}}_{~{}\hat{b}}=d\omega^{\hat{a}}_{~{}\hat{b}}+\omega^{\hat{a}}_{~{}\hat{c}}\wedge\omega^{\hat{c}}_{~{}\hat{b}}.$ From the coefficients of this curvature 2-forms, we read off the curvature components. The result is given by
@@
Through the Einstein equation $G_{ab}=8\pi T_{ab}$, the absence of the off-diagonal components gives $R_{\hat{r}\hat{\theta}}=0$, which leads